THE HYPERFINE SPECTRUM OF RbCl

D. OLSON, J. CEDERBERG, M. WELGE, \underline{M. FEIG}, N. LONDON, T. HÖFT, M. ETTEN, and J. McKINZIE, Physics Department, St. Olaf College, Northfield, MN 55057.

A molecular beam electric resonance spectrometer has been used to observe pure hyperfine transistions of all four isotopic forms of rubidium chloride, with a linewidth of about 150 Hz. So far 179 observed lines have been identified as belonging to ^{85}Rb^{35}Cl, and tentatively fitted to determine the hyperfine constants, including nuclear quadrupole and spin-rotation interactions for both nuclei, and the tensor and scalar spin-spin interactions, for vibrational states v = 0−3 and rotational states J = 1−6. An additional 103 lines have been identified and fitted for^{87}Rb^{35}Cl. The possibilities of a rubidium nuclear octupole interaction, and (in the case of ^{85}Rb) a nuclear electric hexadecapole interaction, are also being considered.